Solve the equation 4x-3 = 2x2 + 5x -6. Please indicate what you graphed and what you got the calculator to do.


1
Expert's answer
2021-11-08T13:17:21-0500

The graph of the quadratic function "y=2x^2+5x-6" is parabola.


"y=2x^2+5x-6"

"=2\bigg(x^2+2(\dfrac{5}{4})x+(\dfrac{5}{4})^2\bigg)-2(\dfrac{5}{4})^2-6"

"=2\big(x+\dfrac{5}{4}\big)^2-\dfrac{73}{8}"

"Vertex:\big(-\dfrac{5}{4}, -\dfrac{73}{8}\big)"

"y" -intersection: "x=0, y=2(0)^2+5(0)-6=-6"

Point "(0, -6)"


"x" -intersection(s): "y=0, 0=2\big(x+\dfrac{5}{4}\big)^2-\dfrac{73}{8}"

"2\big(x+\dfrac{5}{4}\big)^2=\dfrac{73}{8}"

"\big(x+\dfrac{5}{4}\big)^2=\dfrac{73}{16}"

"x+\dfrac{5}{4}=\pm\sqrt{\dfrac{73}{16}}"

"x=-\dfrac{5}{4}\pm\dfrac{\sqrt{73}}{4}"

"x_1=-\dfrac{5}{4}-\dfrac{\sqrt{73}}{4}"

"x_2=-\dfrac{5}{4}+\dfrac{\sqrt{73}}{4}"

Point "\big(\dfrac{-5-\sqrt{73}}{4}, 0\big)," Point "\big(\dfrac{-5+\sqrt{73}}{4}, 0\big)."


"x=1: y=2(1)^2+5(1)-6=1"

Point "(1, 1)."


"x=-2: y=2(-2)^2+5(-2)-6=-8"

Point "(-1, -9)."


The graph of the linear function "y=4x-3" is a straight line.

"x=0: y=4(0)-3=-3" .

Point "(0, -3)"


"y=0: 0=4x-3=>x=\dfrac{3}{4}"

Point "\big(\dfrac{3}{4}, 0\big)"

Points of intersection "(1, 1), (-1.5, -9)"

Check


"2(1)^2+5(1)-6=4(1)-3"

"1=1, True"


"2(-1.5)^2+5(-1.5)-6=4(-1.5)-3"

"-9=-9, True"



Solve the equation algebraically


"2x^2+5x-6=4x-3"

"2x^2+5x-6-4x+3=0"

"2x^2+x-3=0"


"D=(1)^2-4(2)(-3)=25"

"x=\dfrac{-1\pm\sqrt{25}}{2(2)}=\dfrac{-1\pm5}{4}"

"x_1=\dfrac{-1-5}{4}=-1.5"

"y_1=4(-1.5)-3=-9"

"x_2=\dfrac{-1+5}{4}=1"

"y_2=4(1)-3=1"

"x_1=-1.5, x_2=1"